OpenStudy (anonymous):

4y-16+8y=-4 4y-8y ? =-4y -4y-16=-4 ????

OpenStudy (anonymous):

whats the next step o: ?

OpenStudy (anonymous):

ok so you have \[4y-16+8y=-4\] correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

First things first....add the y's on the left hand side of the equal sign \[(4y+8y)-16=-4\] I'm putting parenthesis to point out that I want you to add the y's first and leave the rest alone for now.

OpenStudy (anonymous):

what would you get?

OpenStudy (anonymous):

12y - 16 = -4

OpenStudy (anonymous):

awesome! now how do we get the -16 to the right hand side? make a guess :)

OpenStudy (anonymous):

by subtracting 16 to the other side?

OpenStudy (anonymous):

almost :) you would add 16 take a look... 12y (- 16 +16)= -4+16 so (-16+16) is zero correct?

OpenStudy (anonymous):

then you are left with 12y=-4+16

OpenStudy (anonymous):

I GET IT THAANKSS :))!!

OpenStudy (anonymous):

no prob...so what's the answer?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

yw